by TEI TEI

02/26/08 – Final paperwork, O&M’s and As-Builts have been sent to the County for closeout. This job is complete.

02/21/08 – TTC received the final pay application on this project and signed it immediately and sent it to Tony for closeout on this project. ThermalTech agreed to do the as-built drawings for Tony on this project.

12/05/07 – Today we worked on allowance cost and change order submissions by PHB. TTC asked for finalization and closeout paperwork from Bernie on the job. 3 allowance have to be finalized and TTC asked the County if there were any other tasks on this project they needed done with the remaining allowance money.

11/15/07 – PHB reported today that the equipment went in exactly as drawing except for some minor change in piping right above the heat exchanger. TTC said he would accept a markup on this change and incorporate it into the electronic dwgs for the job and give it to the County for final As-Builts. PHB is working on final closeout documents this week and should have them to TEI for final approval very soon, they only items they are waiting on is JCI sequences.

10/17/07 – JCI does not have the right sensors in the HTX system for their controls. We are investigating replacing them or using an old one. this needs to be resolved quickly.

10/01/07 – TTC & GD visited the site last week and reviewed the project installation. The job was very well done by PHB. GD requested that the strainers be installed in the horizontal position to alleviate water hammer. This is an arduous task for PHB but they are going it today. TTC approved payapp #2. JCI said they were almost finished with controls but had to add flex tight to the controls at the valves. Howard was going to run the system in manual to see how it works. TEI still must provide a punchlist for this project.

09/12/07 – Howard called with a questions about isolating the PRV for maintenance. TTC said it was isolated on the PID. TTC will stop down tomorrow to review with contractors and will setup a progress meeting with GD to go over design.

09/03/07 – In the past week we have approved a pad extension onto the front of the existing HTX pad layout to allow the piping to be connected to the existing piping easier. We have approved new butterfly isolation valves from the allowance monies and TTC approved payapp #01 for the project. In the field TTC approved a slight modification to the layout of the steam traps for easier maintenance and TTC and Walt discussed the atmospheric steam vent that is supposed to be connected to a tube inside the trap and came shipped separately.

08/23/07 – PHB asked about new steam control valves vs old. TEI directed them to the Add01 where hot water control valves are reused but not steam valves. Steam are new CV-27 types and 3″ in size. Howard asked about HTX loads and copper tube sheet strength. TTC asked GD and TZ for info. TTC scheduled a field trip on Friday to the building to review the progress of the project.

08/20/07 – PHB did get the new steam isolation valve installed. The Spence valve and 2″ gate valve was also replaced. Howard said that 2 of the hot water isolation valves are corroded and need to be replaced. TTC asked PHB to give us a price to replace this valves. These particular valves are not more than 5 or 6 years old as they were put in during the

08/17/07 – PHB is installing the steam valve, Spence valve and 2″ gate tonight after hours. TTC will check in with JC tomorrow on progress.

08/03/07 – TTC gave PHB requested information on the Spence valve, isolation valve and existing steam pressures. The Spence valve has unions on both sides for easy replacement. The isolation valve shows 6 inch on the dwgs and with the insulation on the piping, there is no easy way to confirm it isn’t 6″. The steam pressure is 65 psig max for normal operating pressure so PHB’s request to buy a 150 psig isolation valve will be fine. TTC & HS picked the location so that the HTX could be isolated from the system for this work.

05/30/07 – TTC contacted PHB about the equipment orders. The HTX vendor asked for more money for sizing the HTX at 10 psig but TEI emailed PHB and showed that 10 psig was called for on the schedule. This may hold up equipment ordering.

05/21/07 – TEI returned submittals via email to Bernie Tobergte today. Now comes the equipment order and scheduling. TEI will deliver hard copies to Tony and Howard.

05/14/07 – TEI has all submittal data and will review on 05/16/07 and return to PHB by week’s end.

03/14/07 – TM & TTC have contract and are ready for kickoff meeting.

02/01/07 – Contracts have been sent to BOCC for confirmation and then we will kick off the project.

01/02/07 – TEI has submitted all contract and bid documents to Facilities for Prosecutor review but as of this date, we have not had any information back on this project.